OpenStudy (anonymous):

if the length of the hypotenuse of a right triangle is 10, find the length of the other sides when the area is a maximum. ans. 5sqr. 2 x 5sqr. 2

OpenStudy (anonymous):

u doin calculus?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):



OpenStudy (anonymous):

u need to find x such that area is maximim

OpenStudy (anonymous):

Am i going to substitute it atthe eqn 1/2 x (sqrt (100-x^2) ?

OpenStudy (anonymous):

right area\(=f(x)=\frac{1}{2}x \sqrt{100-x^2} \)

OpenStudy (anonymous):

easily u can see 0<x<10 so differentiate f(x) and find the x (solving f'(x)=0 ) for which area is max

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Welcome